OpenStudy (anonymous):

Write the expression 125m^6 - 8n^3 in factored form.

OpenStudy (smileyxl3):

\[(5m^2-2n)(25m^4+10m^2n+4n^2)\]

OpenStudy (anonymous):

could you explain how you got that please?

OpenStudy (smileyxl3):

Use this formula \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

(125m^6)^3-(-8n^3)^3= .. like that?

OpenStudy (smileyxl3):

Well Since both terms are perfect cubes, find the values for a and b.\[a^3=125m^6 so, a=5m^2\] and \[b^3=8n^3, so, b=2n\]

OpenStudy (anonymous):

and plug those into the equation instead?

OpenStudy (smileyxl3):

The binomial can be factored using the difference of cubes formula, because both terms are perfect cubes. So you would do \[(5m^2-2n)(25m^4+10m^2n+4n^2)\]

OpenStudy (anonymous):

okaay

OpenStudy (anonymous):

thanks